OpenStudy (anonymous):

feeling really tired plz help... -5+a=21

OpenStudy (anonymous):

solve for a

OpenStudy (anonymous):

its almost like a free medal

OpenStudy (anonymous):

-5+a=21 +5 +5 a = 26

OpenStudy (anonymous):

ok thx